Part A Analyzing Position-Time Graphs Posstion tom) Lab 01 Motion in One Dimension PHYS 1111L Example Data Tables Data Table for Part A- Analyzing Position-Time Graphs 0-2 sec 2-4 sec 4-6 sec 6-8 sec 8-10 sec
Answer #1

Intenval 30 30-0 = 30 30 60 |66-40=20 6-8 6 20=15 | 1.5+5+10ナ1St5 2- 2. 20-10 2v=(xfinal-xinitial )/time interval.